OpenStudy (anonymous):

Nick is now seven times as old as Harry. In 6 years, Nick will be five times as old as Harry. How old will Harry be in 6 years? 12 16 18

OpenStudy (anonymous):

18 is your answer don't forget to medal :D

OpenStudy (anonymous):

lol thanks

OpenStudy (anonymous):

let harrys present age = x years then nick is 7x years in 6 years harrys age = x+6 nicks age = 7x + 6 ATQ 7x + 6 = 5(x+6) => 7x + 6 = 5x + 30 => 7x - 5x = 30 - 6 => 2x = 24 => x = 24/2 = 12 so Harry's present age = 12 years After 6 years he will be 12+6 = 18 years

OpenStudy (anonymous):

incase you want to know how.

OpenStudy (anonymous):

i did thanks